Please explain each step in the solution

6. Find the equivalent capacitance of the combination. Assume that Ci is 10 uF, C2 is 5uF and C3 is 4uF (2 marks)
Answer #1

we knowo thout Capacitors are connected in series then they gets equally charged +2..-8 +0.-879-8 sa G son replacing all capaw, in parallel Combination, voltage drop across each Capacitor is same. & ♡ Geg - so, we can say that 8=8, +62 +63 Ceq r = cuin parallel I we com see that c, and ly are - Combination to each other, so equivalent of c, and ca. let G be equipalent, thegiven c, zlour g=5.uf za 4MF cea = (4x10-6) (15 x 10-6) 19x10rb So, ceq = 3.15 ME capacitance of the - this is the required e